6. Constant-growth rates One of the most important components of stock valuation is a firm’s estimated growth rate. Financial statements provide the information needed to estimate the growth rate. Kenji Taylor, an equity research analyst at Taylor Advisors, believes in efficient markets. He has been following the mining industry for the past 10 years and needs to determine the constant-growth rate that he should use while valuing Pan Asia Mining Co. Kenji has the following information available: • Pan Asia Mining Co.’s stock (Ticker: PAMC) is trading at 16.25 per share. • The company has forecasted net income and book value of equity for the coming year to be $1,025,700 and $8,027,500, respectively. • The company has also been paying dividends for the past 8 years and has maintained a dividend payout ratio of 32.50%. Based on this information, Kenji’s forecast of PAMC’s growth rate in earnings and dividends should be . Which of the following statements accurately describes the relationship between earnings and dividends, when all other factors are held constant? Long-run earnings growth occurs primarily because firms retain earnings and reinvest them in the business. Dividend growth and earnings growth are unrelated. Paying a higher percentage of earnings as dividends will result in a higher growth rate.
Answer #1

Expected Book Value of Equity = $ 8027500 and Expevted Net Income = $ 1025700

Return on Equity = (1025700 / 8027500) x 100 ~ 12.77 %

Net Income = $ 1027500 and Dividend Payout Ratio = 32.5 %

Retention Ratio = (100 - 32.5) = 67.5 % or 0.675

Sustainable Growth Rate = Retention Ratio x ROE = 0.675 x 12.77 ~ 8.6247 % or 8.62 %

Growth is as is observed above is equal to the product of the retention ratio and the ROE. Hence, growth is dependent directly on the retention ratio which is nothing but the amount of earning that is retained and reinvested in a business. Dividend Growth is directly related to earnings growth and is infact both are equal. Higher dividend payments imply higher dividend payout ratio and lower retention ratio, thereby leading to lower and not higher growth. Hence, statement 1 is the correct option.